OpenStudy (anonymous):

how do you factor (x-4)^5 ?

OpenStudy (anonymous):

That is already the factored form

OpenStudy (anonymous):

If you mean expanding, you can use pascals triangle

OpenStudy (anonymous):

and how do you do that?

OpenStudy (anonymous):

like how do you evaluate it when its (x-4)(x-4)(x-4)(x-4)(x-4) ?

OpenStudy (anonymous):

You can do it the long way, as in multiply it by itself 5 times or use pascals triangle as a shortcut, looking at the link

OpenStudy (anonymous):

Long way is just distribution, (x-4)(x-4) = x^2 -8x +16 (x^2 -8x +16)(x-4) = x^3 .... and so on

OpenStudy (anonymous):

and what is pascals triangle?

OpenStudy (anonymous):

look at the link i sent

OpenStudy (anonymous):

where do i look for that?

OpenStudy (anonymous):

thank you for helping !